OpenStudy (anonymous):

Can anyone help please If you have 5x - 4y = 3 and y = 2x +4, then use substitution.

OpenStudy (anonymous):

5x-4(2x+4)=3 5x-8x-16=3 -3x-16=3 -3x=19 x=-19/3